A box is made without a lid by notching out equal sized squares from the four corners of a poster board and folding up the sides. The poster board measures 18in. X 15in. The length of the side of the notch is 3in. What is the volume of the box?
I thought that I should multiply 18X15X3, but I think this is wrong. I'm not sure why. Can you help? Thanks...John

Answer by checkley77(12844)

The first thing you need to do is draw the dimentioned board, notch out the four corners THEN write the formula.
(18-2*3)(15-2*3)3=volume of the box.
vol=(18-6)(15-6)3
vol=12*9*3
vol=324 in.^3 answer.